Find, showing all working, a recursive de nition of the sequence with general

term

tn = 3 (n + 1)!/2n ; n >= 1


1
Expert's answer
2022-06-06T23:23:42-0400

"t_n =\frac{ 3 (n + 1)!}{2n} ; n \geq1"; will be the general term of the sequence.


Let's find the next term:

"t_{n +1}=\frac{ 3 ((n + 1)+1)!}{2(n+1)}=\frac{ 3 (n+2)!}{2(n+1)}="


"=\frac{ 3 (n+2)(n+1)n\times...\times1}{2(n+1)}=\frac{ 3 (n+2)(n+1)!}{2(n+1)}\frac{n}{n}="


"=\frac{ 3 (n+1)!}{2n}\frac{n(n+2)}{n+1}=t_n\frac{n(n+2)}{n+1}=\frac{n(n+2)}{n+1}t_n."


Also we have to find the first term:

"t_1 =\frac{ 3 (1 + 1)!}{2\cdot 1}=\frac{3\cdot 2!}{2}=\frac{3\cdot 2}{2}=3."


So, the recursive de finition of the sequence is as follows:

"t_1=3,"

"t_{n+1}=\frac{n(n+2)}{n+1}t_n."


Answer:

"t_1=3,"

"t_{n+1}=\frac{n(n+2)}{n+1}t_n."



But if we have "2^n" in denominator, the solution is the next:

"t_n =\frac{ 3 (n + 1)!}{2^n} ; n \geq 1"; will be the general term of the sequence.

Let's find the next term:

"t_{n +1}=\frac{ 3 ((n + 1)+1)!}{2^{n+1}}=\frac{ 3 (n+2)!}{2^{n+1}}="


"=\frac{ 3 (n+2)(n+1)n\times...\times1}{2^n\cdot 2^1}=\frac{ 3(n+1)!}{2^n}\frac{n+2}{2}=t_n\frac{n+2}{2}."


Also we have to find the first term:

"t_1 =\frac{ 3 (1 + 1)!}{2^1}=\frac{3\cdot 2!}{2}=\frac{3\cdot 2}{2}=3."


So, the recursive defi nition of the sequence is as follows:

"t_1=3,"

"t_{n+1}=\frac{n+2}{2}t_n."


Answer:

"t_1=3,"

"t_{n+1}=\frac{n+2}{2}t_n."